TMS Meeting Minutes, May 13, 2021

Attendees: Andrew, Breann, Heejoo, Matthieu, Olga, Trenton, Yang (Zoom Meeting)

Etalon Work on Etalon Software:
  • Etalon has been working on fixing their software, turns out it was a Windows driver error
  • The driver to control the laser diode (3rd party software) is the culprit, where the driver times out when running for long periods
  • Etalon is able to reproduce the issue on their side
  • Etalon is trying a solution involving immediate reconnection during a closed port fault from the driver
  • Etalon still has more testing to run, and Yang and Heejoo will be assisting their team with this using TCP/IP
  • Primarily, they want to have confirmation that the Etalon updates work in conjunction with the updated TCP/IP code for bino mode
  • Plan is to run the system for several hours, obtain a failure (naturally) which will be logged, and then show that the updated Etalon code will recover from a failure gracefully
  • This test will probably take place tomorrow, pending confirmation from Mark at Etalon
Yang Finished the TMS Bino Code, Planning Closed-Dome Test
  • Planning to run basic tests next week to check basic functionality of the system, the channel numbering, etc.
  • Heejoo will ask LeRoy to reconnect all fibers following Etalon's testing
  • Heejoo has finished the post-processing of the data tracker data to give
  • Andrew proposes developing a TMS lookup table for both sides to keep collimation and pointing
    • Move telescope up and down in elevation, sense the deviations of the relative positioning
    • Similarly with copointing, by sensing relative offsets of DX and SX
    • A refined LUT may be developed for TMS use as an improvement based on direct measurements
    • Also a really nice workout for the software
    • Could also be scripted, s.t. take a few measurements, change elevation, few measurements, etc. ... resolve changes, and more
    • This could even be done next week, pending a successful test tomorrow
  • Clarification on collimation vs. co-pointing
    • Without being on-sky, you can't do true co-pointing, only primary mirror relative to hub
    • If differential motion from DX relative to SX, can't sense that
    • Assuming common deflections of the primary mirror relative to the hub, able to decompose into two primary components
      • Component of change inducing on-axis optical aberrations
      • Component of change changing pointing without introducing aberrations
    • Ensure that the pointing on both sides keeps you locked to the hub frame, without introducing aberration (coma-neutral rotation vs. pointing neutral rotations)
Solid Retro Delivered to GMTO
  • Breann has confirmed that they have received the solid retro in Pasadena
Upcoming Testing Opportunities for On-Sky Passive Mode
  • First time in two years, binocular TMS will be ready to go on-sky
  • It looks like there is some availability on the 30th
  • Andrew is continuing to prepare the commissioning document and will continue working on this
  • Andrew will try to get the closed-dome commissioning activities done first
  • This commissioning document could be a good background for observers on what to expect when running TMS
  • Olga suggests that we should try to get some on-sky time soon, once we are ready
    • To go on-sky, we need to prove that TMS doesn't make things worse, that collimation doesn't diverge, that TMS doesn't diverge
    • Shared risk is mitigated by the fact that the observer will be able to watch the data as it comes in, and is able to stop and correct the system if it starts to fall apart
    • There are specific groups which might choose to use LBC when they wouldn't have before, because of the existence and performance gains achieved by the use of TMS
  • If we expect to have a night, demonstrating overall TMS performance, while able to fail gracefully, Andrew proposes that we should be able to propose that the system be used by observers
  • Olga planning to prepare a document for observers
Problem of Non-planarity of Chips + Tilt of Entire Focal Plane
  • The focal plane shims are not of the proper thickness, so the overall focal plane is tilted
  • Olga planning to do some testing before the observatory goes offline to characterize the tilts of the focal plane

  • Andrew requests a power consumption measurent during laser + Etalon operation, both in the static case and while running in a loop
  • Olga has a kill-a-watt at her home which she has offered up for use with this power consumption testing

-- TrentonBrendel - 13 May 2021
Topic revision: r1 - 14 May 2021, TrentonBrendel
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback